Write each fraction in simplest form:
17/51

42/55

Answer by CharlesG2(834)

Write each fraction in simplest form:
17/51
42/55
17/51 = 1*17 / 3*17 = 1/3
42/55 = 2 * 3 * 7 / 5 * 11 that can not be put simpler so it stays 42/55